Can someone help me solve this problem?
I have a right triangle.
On the left side (the verticle left side ) is 7
and on top it is 25
I need to find the missing length on the botton for a right triangle.
Can you explain how you come up with the missing length in detail?
Thanks,
Sher

Answer by venugopalramana(3286)

Can someone help me solve this problem?
I have a right triangle.
On the left side (the verticle left side ) is 7
and on top it is 25
I need to find the missing length on the botton for a right triangle.
Can you explain how you come up with the missing length in detail?
Thanks,
Sher
IT IS GIVEN BY
HYPOTENUSE^2=SIDE^2+ANOTHER SIDE^2
25^2=7^2+SIDE^2
SIDE^2=625-49=576
SIDE=24